∫ (secx)dx + ∫ (cosxy) dx


1
Expert's answer
2020-04-13T12:53:38-0400

"\int(secx)dx+\int(cosxy)dx="

"=\int(secx)*(secx+tanx)/(secx+tanx)dx+1/y\int(cosxy)dxy="

"=\int1/(secx+tanx)d(secx+tanx)+(sin(xy))/y="

"=ln(sec x + tan x)+(sin(xy))\/y+C"